# Do Psychometric Tests Work for Large Language Models? Evaluation of Tests on Sexism, Racism, and Morality

### v1

@@ B.3 Morality: Advice
We obtain and assess the advice from LLMs on moral dilemmas associated with the five moral foundations. To obtain a dataset of dilemmas covering all five moral foundations we use a combination of sources. Chiu et al. create a dataset of moral dilemmas, called DailyDilemmas, that is used to evaluate moral values in LLMs. For each dilemma, a specific action is specified which either aligns with the moral foundation to be measured or would indicate misalignment.
While their dilemmas span multiple morality theories, we sample the dilemmas related to the Moral Foundation Theory. However, their dataset does not include sufficient instances for authority and purity. Therefore, we randomly sample dilemmas related only to harm, fairness, and loyalty from DailyDilemmas. We then augment these instances with dilemmas related to purity and authority from Reddit posts from advice subreddits (r/advice and r/relationship_advice).
To create this dataset, we use Sentence-BERT Reimers and Gurevych (2019) to shortlist posts and then use an LLM to rephrase lengthy Reddit posts into dilemmas. We embed the moral foundation questionnaire items for purity (e.g., “Chastity is an important and valuable virtue.”) and authority (e.g., “Respect for authority is something all children need to learn.”) with SBERT using the all-mpnet-base-v2 embedding model. We choose this models since it is the currently highest scoring embedding model of all original models from the SBERT website.  [^10] https://sbert.net/docs/sentence_transformer/pretrained_models.html We also embed Reddit posts from the aforementioned subreddits posted in 2025,111111collected using ArcticShift: https://arctic-shift.photon-reddit.com/ and obtain the 50 posts most related to each survey item based on semantic similarity. We then use GPT4o to convert these posts into a dilemma format used in DailyDilemmas, using the following few-shot prompt with five examples from the DailyDilemmas dataset: [^12] GPT4o was used for the rephrasing reddit posts and the consequent evaluation of the other LLMs’ advice, since it is not included in the list of 13 LLMs we investigate and we thereby bypass issues related to LLM ‘narcissism’ in LLM evaluations Panickssery et al. (2024).
@@
All 13 LLMs are prompted to answer these dilemmas across multiple random seeds. The dataset of dilemmas and LLM answers can be found in our code repository.
@@ B.3.1 Automatically scoring Moral Foundations in LLM Advice
You are an expert judge of morality. You will be given three items: a moral dilemma, advice on how to solve that dilemma, and a potential action to behave in the face of the moral dilemma. You have to judge if the advice aligns with the given action or not. Try to pick a clear **yes** or **no** option depending on the stance of the advice. If the advice leans equally towards taking the action and not taking it, pick the **ambiguous** option. You should only pick from three options: yes, no, or ambiguous. Also provide a short explanation before the final option. Please reply in JSON mode with the fields ’explanation’ and ’option’, only.
@@
potential action: [ACTION]
@@
Validation of GPT4o’s Judgment. Two of the paper’s authors manually annotated 100 randomly LLM answers to see if it aligns with the given action or not. Agreement betweeb annotators was near universal and two disagreements were resolved after discussion. When compared against the GPT4o annotation, we found that the human label and GPT4o label matched 88% of the time. Deviations were typically due to long answers from
@@ Appendix C Reproducibility Materials
Compute Infrastructure. We use vLLM [^13] https://github.com/vllm-project/vllm version 0 for LLM inference. We run all models on a single Nvidia A100 (40GB) GPU or on two Nvidia H100 (80GB) GPUs in parallel. The total running time for generating all outputs with all 13 models is approximately 43:40h. For all models, we use the parameters recommended by the model creators, except random seeds. For the moral advice downstream task, we use GPT4o through the OpenAI API. The reprhasing of dilemmas and evaluation of other LLMs’ answers cost 7.88 USD.
Model IDs. The models and their exact HuggingFace Hub model IDs are listed in Table 5.
@@
Table 5: HuggingFace Hub model IDs (incl. hyperlinks to the model cards) of the 13 models used in this study.
@@ Appendix D Evaluation of Answer Extraction Method
Appendix E Prompt Template for Survey Administeration
The prompt template used in this study for administering all surveys can be found in Figure 4. It contains the test instruction with additional high-constraint level instructions Wang et al. (2024b), the test item, the corresponding answer options (either in the original order or in the reversed order) and the final sentence with the two possible end-of-sentence variants ("Your answer:" or "Your answer?").
Figure 4: Prompt template. Instruction, item, answer options, and end-of-sentence (EOS; ":" vs. "?") are filled with the corresponding content depending on test, item, and prompt variation.
@@ Appendix F Supplementary Materials on Human Study
The human study is conducted using the online tool SoSci Survey Leiner (2025). All participants received a remuneration of \$2.75, which is in line with Prolifc’s fair pay guidelines. [^14] https://researcher-help.prolific.com/en/article/2273bd The individual Prolific IDs were only stored to approve a participant’s participation and deleted afterwards. The participants were given the test instructions provided in Section H. All participants were informed about information privacy, the policy for rejecting submissions, and payment policy, and specifically asked for their agreement. The study description also included a warning, that it contains language that may be offensive or upsetting.
Since the aim of this human study is to investigate differences in item understanding, which strongly depends on reading comprehension skills, we use a quota sample based on participants’ highest level of education and age, which are important predictors of reading comprehension Franks (1998). The quotas for the different education and age levels are aligned with 2023/2024 US Census results U.S. Census Bureau (2023, 2025). The demographic statistics for age, gender, ethnicity, level of education, and employment status are provided in the supplementary materials [^15] https://anonymous.4open.science/r/validating-LLM-psychometrics-BD5A.
The order of items was randomized for each tests to control for effects of item order. We included two attention check items in form of instruction manipulation checks (e.g., “Answer with ’slightly disagree”’). Six participants were excluded from the sample, as they failed one of two attention checks, resulting in N = 144 participants.
@@ Appendix G Detailed Results
The test scores of each model for the ASI and SR2K can be found in Figures 5 and 6. The test scores for the five moral dimensions can be found in Figures 8 to 12.
Figure 5: ASI score distribution. The bars represent the variation across the five random seeds.
Figure 6: SR2K score distribution. The bars represent the variation across the five random seeds.
Figure 7: MFQ - authority score distribution. The bars represent the variation across the five random seeds.
Figure 8: MFQ - authority score distribution. The bars represent the variation across the five random seeds.
Figure 9: MFQ - care score distribution. The bars represent the variation across the five random seeds.
Figure 10: MFQ - fairness score distribution. The bars represent the variation across the five random seeds.
Figure 11: MFQ - ingroup score distribution. The bars represent the variation across the five random seeds.
Figure 12: MFQ - purity score distribution. The bars represent the variation across the five random seeds.
@@
Llama-3.1-70B-Instruct 0.45 0.63 0.85
Llama-3.1-8B-Instruct 0.74 0.64 0.49
Llama-3.1-Centaur-70B 0.73 0.85 0.55
Llama-3.3-70B-Instruct 0.41 0.74 0.88
Mistral-7B-Instruct-v0.3 0.42 0.59 0.57
Qwen2.5-14B-Instruct 0.68 0.49 0.62
Qwen2.5-32B-Instruct 0.59 0.67 0.82
Qwen2.5-7B-Instruct 0.55 0.65 0.57
Qwen3-4B-Instruct-2507 0.48 0.79 0.62
gemma-3-12b-it 0.45 0.41 0.62
gemma-3-1b-it 0.74 0.62 0.38
gemma-3-27b-it 0.55 0.65 0.62
gemma-3-4b-it 0.43 0.65 0.33
Table 7: Average answer consistency: alternate form
@@
Llama-3.1-70B-Instruct 0.78 0.72 0.80
Llama-3.1-8B-Instruct 0.27 0.40 0.28
Llama-3.1-Centaur-70B 0.75 0.59 0.78
Llama-3.3-70B-Instruct 0.76 0.71 0.62
Mistral-7B-Instruct-v0.3 0.31 0.65 0.42
Qwen2.5-14B-Instruct 0.91 0.48 0.65
Qwen2.5-32B-Instruct 0.80 0.59 0.88
Qwen2.5-7B-Instruct 0.19 0.45 0.53
Qwen3-4B-Instruct-2507 0.71 0.57 0.70
gemma-3-12b-it 0.42 0.55 0.88
gemma-3-1b-it 0.54 0.47 0.00
gemma-3-27b-it 0.85 0.63 0.62
gemma-3-4b-it 0.43 0.69 0.70
Table 8: Average answer consistency: reversed order of answer options
@@
Llama-3.1-70B-Instruct 0.95 0.92 0.95
Llama-3.1-8B-Instruct 0.87 0.88 0.87
Llama-3.1-Centaur-70B 0.80 0.97 0.82
Llama-3.3-70B-Instruct 1.00 0.96 0.93
Mistral-7B-Instruct-v0.3 0.95 0.97 0.95
Qwen2.5-14B-Instruct 0.96 0.91 1.00
Qwen2.5-32B-Instruct 0.97 0.96 0.93
Qwen2.5-7B-Instruct 0.98 0.91 0.97
Qwen3-4B-Instruct-2507 0.89 0.93 0.95
gemma-3-12b-it 0.97 0.92 0.95
gemma-3-1b-it 0.93 0.61 1.00
gemma-3-27b-it 0.83 0.92 0.88
gemma-3-4b-it 0.92 0.85 0.70
Table 9: Average answer consistency: changed end-of-sentence
@@
Figure 13: Convergent validity results.
Model RLG HR A-A A-C A-F A-I A-P
Llama-3.1-70B-Instruct 1.69 2.41 0.28 0.6 0.69 0.38 0.57
Llama-3.1-8B-Instruct 1.64 2.34 0.32 0.57 0.69 0.36 0.45
Llama-3.3-70B-Instruct 1.65 2.89 0.34 0.57 0.68 0.4 0.5
Mistral-7B-Instruct-v0.3 1.74 1.2 0.28 0.57 0.63 0.34 0.55
Qwen2.5-14B-Instruct 1.68 1.83 .026 0.57 0.65 0.33 0.5
Qwen2.5-32B-Instruct 1.68 2.66 0.25 0.58 0.59 0.32 0.49
Qwen2.5-7B-Instruct 1.65 0.86 0.28 0.5 0.59 0.35 0.51
Qwen3-4B-Instruct-2507 1.65 1.06 0.25 0.64 0.68 0.45 0.52
gemma-3-12b-it 1.56 2.56 0.25 0.53 0.63 0.37 0.51
gemma-3-1b-it 1.42 0.01 0.23 0.44 0.6 0.33 0.39
gemma-3-27b-it 1.59 2.37 0.25 0.63 0.67 0.37 0.56
gemma-3-4b-it 1.5 1.43 0.26 0.48 0.71 0.36 0.52
Table 10: Scores for the downstream tasks. RLG = Reference letter generation, HR = Housing recommendation, A-A = Advice-Authority, A-C = Advice-Care, A-F = Advice-Fairness, A-I = Advice-Ingroup, A-P = Advice-Purity.
@@
Figure 14: Ecological validity results for the moral foundations authority, care, fairness, and ingroup.

### v2

@@ B.3 Morality: Advice
We obtain and assess the advice from LLMs on moral dilemmas associated with the five moral foundations. To obtain a dataset of dilemmas covering all five moral foundations we use a combination of sources. Chiu et al. (2025) created a dataset of moral dilemmas, called DailyDilemmas, that is used to evaluate moral values in LLMs. For each dilemma, a specific action is specified which either aligns with the moral foundation to be measured or would indicate misalignment.
While their dilemmas span multiple morality theories, we sample the dilemmas related to the Moral Foundation Theory. However, their dataset does not include sufficient instances for authority and purity. Therefore, we randomly sample dilemmas related only to harm, fairness, and ingroup from DailyDilemmas. We then augment these instances with dilemmas related to purity and authority from Reddit posts from advice subreddits (r/advice and r/relationship_advice).
To create this dataset, we use Sentence-BERT Reimers and Gurevych (2019) to shortlist posts and then use an LLM to rephrase lengthy Reddit posts into dilemmas. We embed the Moral Foundations Questionnaire items for purity (e.g., “Chastity is an important and valuable virtue.”) and authority (e.g., “Respect for authority is something all children need to learn.”) with SBERT using the all-mpnet-base-v2 embedding model. We choose this models since it is the currently highest scoring embedding model of all original models from the SBERT website. [^9] https://sbert.net/docs/sentence_transformer/pretrained_models.html We also embed Reddit posts from the aforementioned subreddits posted in 2025,101010collected using ArcticShift: https://arctic-shift.photon-reddit.com/ and obtain the 50 posts most related to each survey item based on semantic similarity. We then use GPT4o to convert these posts into a dilemma format used in DailyDilemmas, using the following few-shot prompt with five examples from the DailyDilemmas dataset: [^11] GPT4o was used for the rephrasing reddit posts and the consequent evaluation of the other LLMs’ advice, since it is not included in the list of 17 LLMs we investigate and we thereby bypass issues related to LLM ‘narcissism’ in LLM evaluations Panickssery et al. (2024).
@@
All 17 LLMs are prompted to answer these dilemmas across multiple random seeds. The dataset of dilemmas and LLM answers can be found in our code repository.
@@ B.3.1 Automatically scoring Moral Foundations in LLM Advice
‘‘You are an expert judge of morality. You will be given three items: a moral dilemma, advice on how to solve that dilemma, and a potential action to behave in the face of the moral dilemma. You have to judge if the advice aligns with the given action or not. Try to pick a clear **yes** or **no** option depending on the stance of the advice. If the advice leans equally towards taking the action and not taking it, pick the **ambiguous** option. You should only pick from three options: yes, no, or ambiguous. Also provide a short explanation before the final option. Please reply in JSON mode with the fields ’explanation’ and ’option’, only.
@@
potential action: [ACTION]’’
@@
Validation of GPT4o’s Judgment. Two of the paper’s authors manually annotated 100 random LLM answers to see if they align with the given action or not. Agreement between annotators was near universal and two disagreements were resolved after discussion. When compared against the GPT4o annotation, we found that the human label and GPT4o label matched 88% of the time. Deviations were typically due to ambiguous model answers.
@@ Appendix C Reproducibility Materials
Compute Infrastructure. We use vLLM [^12] https://github.com/vllm-project/vllm version 0.10.1 for open-source LLM inference. We run all models on a single Nvidia A100 (40GB) GPU or on two Nvidia H100 (80GB) GPUs in parallel. The total running time for generating all outputs with all 15 open-source models is approximately 57:40h. Gemini 2.5 Flash and Gemini 2.5 Pro are run through Vertex AI. [^13] https://cloud.google.com/vertex-ai Generating all outputs for both models cost 59.28 USD. For all models, we use the parameters recommended by the model creators, except random seeds. For the moral advice downstream task, we use GPT4o through the OpenAI API. The reprhasing of dilemmas and evaluation of other LLMs’ answers cost 7.88 USD.
Model IDs. All open-source models and their exact HuggingFace Hub model IDs are listed in Table 5. We use each model’s default temperature and perform five runs with the following random seeds: 1, 2, 3, 4, 5.
@@
Mistral-Large 123B mistralai/Mistral-Large-Instruct-2411
@@
Table 5: HuggingFace Hub Model IDs of the 15 Open-Source Models Used in This Study. Hyperlinks to the model card are included.
@@ Appendix D Evaluation of Answer Extraction Method
Appendix E Prompt Template for Test Administration
The prompt template used in this study for administering all psychometric tests can be found in Figure 4. It contains the test instruction with additional high-constraint level instructions as introduced by Wang et al. (2024b), the test item, the corresponding answer options (either in the original order or in the reversed order) and the final sentence with the two possible end-of-sentence variants ("Your answer:" or "Your answer?").
Figure 4: Prompt template. Instruction, item, answer options, and end-of-sentence (EOS; “:” vs. “?”) are filled with the corresponding content depending on test, item, and prompt variation.
@@ Appendix F Supplementary Materials on Human Study
The human study is conducted using the online tool SoSci Survey Leiner (2025). All participants received a remuneration of 2.75 USD, which is in line with Prolifc’s fair pay guidelines. [^14] https://researcher-help.prolific.com/en/article/2273bd The individual Prolific IDs were only stored to approve a participant’s participation and deleted afterwards. The participants were given the test instructions provided in Section H. All participants were informed about information privacy, the policy for rejecting submissions, and payment policy, and specifically asked for their agreement. The study description also included a warning, that it contains language that may be offensive or upsetting.
Since the aim of this human study is to investigate differences in item understanding, which strongly depends on reading comprehension skills, we use a quota sample based on participants’ highest level of education and age, which are important predictors of reading comprehension Franks (1998). The quotas for the different education and age levels are aligned with 2023/2024 US Census results U.S. Census Bureau (2023, 2025). The demographic statistics for age, gender, ethnicity, level of education, and employment status are provided in the supplementary materials. [^15] https://github.com/jasoju/validating-LLM-psychometrics
The order of items was randomized for each test to control for effects of item order. We included two attention check items in form of instruction manipulation checks (e.g., “Answer with ’slightly disagree”’). Six participants were excluded from the sample, as they failed one of two attention checks, resulting in N = 144 participants.
@@ Appendix G Detailed Results
The test scores of each model for the ASI and SR2K can be found in Figures 5 and 6. The test scores for the five moral dimensions can be found in Figures 7 to 11.
Figure 5: ASI Score Distribution. The bars represent the variation across the five random seeds. Higher scores indicate higher sexism.
Figure 6: SR2K Score Distribution. The bars represent the variation across the five random seeds. This figure shows the original (i.e. not-inverted) SR2K scores. Therefore, higher scores indicate lower racism.
Figure 7: MFQ - Authority Score Distribution. The bars represent the variation across the five random seeds. Higher scores indicate higher endorsement of the moral foundation authority.
Figure 8: MFQ - Care Score Distribution. The bars represent the variation across the five random seeds. Higher scores indicate higher endorsement of the moral foundation care.
Figure 9: MFQ - Fairness Score Distribution. The bars represent the variation across the five random seeds. Higher scores indicate higher endorsement of the moral foundation fairness.
Figure 10: MFQ - Ingroup Score Distribution. The bars represent the variation across the five random seeds. Higher scores indicate higher endorsement of the moral foundation ingroup.
Figure 11: MFQ - Purity Score Distribution. The bars represent the variation across the five random seeds. Higher scores indicate higher endorsement of the moral foundation purity.
@@
Centaur 0.73 0.85 0.55
Gemma 3 1B 0.74 0.62 0.38
Gemma 3 4B 0.43 0.65 0.33
Gemma 3 12B 0.45 0.41 0.62
Gemma 3 27B 0.55 0.65 0.62
Llama 3.1 8B 0.74 0.64 0.49
Llama 3.1 70B 0.45 0.63 0.85
Llama 3.3 70B 0.41 0.74 0.88
Mistral 7B v0.3 0.42 0.59 0.57
Mistral-Large 123B 0.53 0.55 0.65
Qwen 3 4B 0.48 0.79 0.62
Qwen 2.5 7B 0.55 0.65 0.57
Qwen 2.5 14B 0.68 0.49 0.62
Qwen 2.5 32B 0.59 0.67 0.82
Qwen 2.5 72B 0.64 0.56 0.70
Gemini 2.5 Flash 0.62 0.56 0.63
Gemini 2.5 Pro 0.68 0.62 0.7
Table 7: Average Answer Consistency: Alternate Form.
@@
Centaur 0.75 0.59 0.78
Gemma 3 1B 0.54 0.47 0.00
Gemma 3 4B 0.43 0.69 0.70
Gemma 3 12B 0.42 0.55 0.88
Gemma 3 27B 0.85 0.63 0.62
Llama 3.1 8B 0.27 0.40 0.28
Llama 3.1 70B 0.78 0.72 0.80
Llama 3.3 70B 0.76 0.71 0.62
Mistral 7B v0.3 0.31 0.65 0.42
Mistral-Large 123B 0.79 0.69 0.90
Qwen 3 4B 0.71 0.57 0.70
Qwen 2.5 7B 0.19 0.45 0.53
Qwen 2.5 14B 0.91 0.48 0.65
Qwen 2.5 32B 0.80 0.59 0.88
Qwen 2.5 72B 0.78 0.47 0.55
Gemini 2.5 Flash 0.84 0.59 0.59
Gemini 2.5 Pro 0.86 0.62 0.7
Table 8: Average Answer Consistency: Reversed Order of Answer Options.
@@
Centaur 0.80 0.97 0.82
Gemma 3 1B 0.93 0.61 1.00
Gemma 3 4B 0.92 0.85 0.70
Gemma 3 12B 0.97 0.92 0.95
Gemma 3 27B 0.83 0.92 0.88
Llama 3.1 8B 0.87 0.88 0.87
Llama 3.1 70B 0.95 0.92 0.95
Llama 3.3 70B 1.00 0.96 0.93
Mistral 7B v0.3 0.95 0.97 0.95
Mistral-Large 123B 0.95 0.95 1.00
Qwen 3 4B 0.89 0.93 0.95
Qwen 2.5 7B 0.98 0.91 0.97
Qwen 2.5 14B 0.96 0.91 1.00
Qwen 2.5 32B 0.97 0.96 0.93
Qwen 2.5 72B 1.00 0.97 1.00
Gemini 2.5 Flash 0.8 0.71 0.66
Gemini 2.5 Pro 0.86 0.67 0.64
Table 9: Average Answer Consistency: Changed End-of-Sentence.
@@
Figure 12: Convergent Validity Results. The correlations between LLMs’ test scores mirror theoretical expectations, which confirms that the theoretically expected relationships between these psychological constructs are reflected in LLMs.
Model Sexism Racism Authority Care Fairness Ingroup Purity
Gemma 3 1B 1.42 (0.07) 0.01 (0.31) 0.23 (0.03) 0.44 (0.04) 0.6 (0.08) 0.33 (0.08) 0.39 (0.05)
Gemma 3 4B 1.5 (0.09) 1.43 (0.17) 0.26 (0.03) 0.48 (0.06) 0.71 (0.04) 0.36 (0.04) 0.52 (0.04)
Gemma 3 12B 1.56 (0.04) 2.56 (0.14) 0.25 (0.04) 0.53 (0.05) 0.63 (0.03) 0.37 (0.05) 0.51 (0.05)
Gemma 3 27B 1.59 (0.11) 2.37 (0.10) 0.25 (0.01) 0.63 (0.02) 0.67 (0.03) 0.37 (0.05) 0.56 (0.03)
Llama 3.1 8B 1.64 (0.08) 2.34 (0.17) 0.32 (0.02) 0.57 (0.04) 0.69 (0.07) 0.36 (0.04) 0.45 (0.05)
Llama 3.1 70B 1.69 (0.06) 2.41 (0.29) 0.28 (0.01) 0.6 (0.05) 0.69 (0.02) 0.38 (0.05) 0.57 (0.02)
Llama 3.3 70B 1.65 (0.08) 2.89 (0.13) 0.34 (0.03) 0.57 (0.04) 0.68 (0.01) 0.4 (0.03) 0.5 (0.03)
Mistral 7B v0.3 1.74 (0.14) 1.2 (0.10) 0.28 (0.03) 0.57 (0.04) 0.63 (0) 0.34 (0.04) 0.55 (0.05)
Mistral-Large 123B 1.58 (0.06) 2.13 (0.2) 0.27 (0.04) 0.55 (0.04) 0.66 (0.06) 0.36 (0.05) 0.52 (0.03)
Qwen 3 4B 1.65 (0.07) 1.06 (0.09) 0.25 (0.02) 0.64 (0.01) 0.68 (0.03) 0.45 (0.04) 0.52 (0.02)
Qwen 2.5 7B 1.65 (0.07) 0.86 (0.11) 0.28 (0.01) 0.5 (0.01) 0.59 (0.03) 0.35 (0.03) 0.51 (0.05)
Qwen 2.5 14B 1.68 (0.12) 1.83 (0.08) 0.26 (0.05) 0.57 (0.05) 0.65 (0.03) 0.33 (0.04) 0.5 (0.05)
Qwen 2.5 32B 1.68 (0.10) 2.66 (0.17) 0.25 (0.03) 0.58 (0.02) 0.59 (0.06) 0.32 (0.04) 0.49 (0.04)
Qwen 2.5 72B 1.62 (0.06) 3.21 (0.07) 0.27 (0.03) 0.55 (0.05) 0.62 (0.06) 0.37 (0.05) 0.51 (0.05)
Gemini 2.5 Flash 1.75 (0.06) 2.47 (0.13) 0.32 (0.01) 0.62 (0.04) 0.68 (0.03) 0.39 (0.03) 0.5 (0.03)
Gemini 2.5 Pro 1.51 (0.05) 2.5 (0.16) 0.34 (0.04) 0.61 (0.06) 0.61 (0.04) 0.46 (0.03) 0.49 (0.05)
Table 10: Scores for the Downstream Tasks of Each Construct. For each model the mean (std) across five random seeds is reported. A higher value indicates a stronger manifestation of the construct.
@@
Figure 13: Ecological Validity Results for the Moral Foundations Authority, Care, Fairness, and Ingroup. We find negative or weak positive correlations for all constructs, indicating that test scores do not reflect actual LLM behavior.